# Advancements and future expansions of the Caribou DAQ system

### v1

@@ 1 Introduction
The development of new silicon pixel detectors for high-energy physics experiments requires versatile and reusable data acquisition (DAQ) systems capable of supporting rapid prototyping, laboratory qualification, and test-beam operation. The Caribou system [1, 2, 3, 4, 5, 6, 7] was developed to meet these needs within several collaborative frameworks — CERN EP R&D, DRD3, AIDAinnova, and Tangerine — and is now used by more than fourteen institutes for the characterization of a wide range of detector prototypes. It combines a custom Control and Readout (CaR) board [8] with a Xilinx Zynq System-on-Chip (SoC) platform running a common firmware and software stack shared across projects. Its architecture emphasizes modularity, reusability, and ease of integration, enabling more than fifteen different detector prototypes to be tested within a unified environment. The CaR board provides key interfaces such as programmable power supplies, voltage and current references, high-speed ADCs, injection pulsers, and configurable I/O lines for detector control and readout. It connects to the SoC through FMC and to detector-specific chip boards via a SEARAY connector, allowing adaptation to various detector designs while maintaining a common DAQ backbone. On the software and firmware side, Caribou integrates two central components via its PetaLinux image builder [9] : Peary [10], a C++ framework offering hardware abstraction, configuration management, and multi-device control via command-line and Python interfaces; and Boreal [11, 12], an FPGA firmware framework providing reusable IP cores and automated synthesis and simulation workflows. The use of Vivado Block Design facilitates the coexistence of shared control infrastructure and user-specific logic, while a strict separation between common and detector-specific components ensures maintainability and scalability. Figure 1 shows a block diagram of the Caribou system hardware architecture. It outlines the key resources provided by the Zynq SoC platform, the CaR board, and the chip board, as well as there interconnections and corresponding signal types.
Figure 1: Block diagram of the Caribou hardware architecture
@@ 2.1 Hardware Design and Validation
To validate the functionality of the CaR board v1.5, a dedicated validation setup was developed. Figure 2(a) shows a picture of the setup, which comprises a Zynq board, an adapter board, the CaR board under test, and a custom breakout board. The Zynq board executes the software and firmware routines used to verify the different functionalities of the CaR board. The adapter board serves as an I/O expander, providing access to all CaR board interfaces by utilizing both FMC connectors on the Zynq board. Finally, the custom breakout board enables loopback routing of LVDS signals, includes SMA connectors for high-speed transceiver lines, and provides voltage and current inputs and outputs for both the slow and fast ADCs. The primary objective of this setup is to validate all functionalities of the CaR board by testing its individual components and I/Os upon the arrival of newly produced boards and before their distribution to users. The currently available routines enable verification of the LVDS lines, high-speed transceiver links, power supplies, and ADCs, and are developed within the Boreal [11] and Peary [10] frameworks.
(a) CaR board v1.5 validation setup
(b) CaR board v2 test board
Figure 2: Pictures of the CaR board v1.5 validation setup (a) and the v2.0 test board (b)
In preparation for the design of the CaR board v2, a simplified test board as shown in Figure 2(b) was developed. This $16.5\times 17.5\ \mathrm{cm^{2}}$ compact board does not include a System-on-Module, and instead implements only a subset of the analog and digital circuitry foreseen for the final design. The test board integrates three adjustable power supplies - one providing up to +15V+15V at $1A$, one down to 15V-15V at $1A$, and one up to +5V+5V at $2A$ — as well as DACs, monitoring ADCs, and a programmable bias source with a $\pm 15V$ range and $\pm 10mA$ current limit. It also includes a DAC with $\pm 15V$ range and $\pm 55mA$ per channel, load circuits for PSU testing, and a programmable overcurrent protection system. All components on the board are accessible and controlled through a single USB-to-I2C/SPI interface. The purpose of this test board is to validate the performance of the power supplies, analog circuits, and peripheral components such as ADCs and DACs. It serves as a platform to evaluate different design options, including integrated DC-DC converter modules, discrete regulators, and high-/low-current source configurations. The targeted design improvements for the CaR board v2 include extended voltage and current ranges for the power supplies and current sources, the addition of a negative voltage supply, and enhanced overcurrent protection implemented on an FPGA.
CaR board v2 test boards have been produced and are fully operational, with extensive validation currently in progress. The current and bias sources have been tested, while power supply testing is ongoing. All tested current sources (two per board on two boards) show a gain error performance within $\pm 0.1-0.25\%$ of the full-scale range. The power supplies are functional, and detailed performance results will be reported as testing continues.
@@ 2.3 Embedded Software Architecture Design Revision
Figure 3: Simplified UML diagram of the revised Peary architecture
@@ 3 Conclusion
The developments presented in this paper mark a significant step forward in the evolution of the Caribou system. On the hardware side, the validation of the CaR board v1.5 and the design of the CaR board v2 test board have established a robust foundation for the upcoming CaR board v2. These efforts have enabled early testing and optimization of key analog and power-supply circuits, ensuring improved reliability and performance in the next generation of the platform. In parallel, the complete redesign of the Peary embedded software has introduced a modular and extensible architecture that can support multiple Zynq platforms and future hardware revisions. The introduction of the Hardware Abstraction Layer and the concept of Resources provides a flexible framework for integrating new boards and detector devices with minimal reconfiguration. Together, these advancements considerably strengthen the modularity, scalability, and maintainability of the Caribou system.

### v2

@@ 1 Introduction
The development of new silicon pixel detectors for high-energy physics experiments requires versatile and reusable data acquisition (DAQ) systems capable of supporting rapid prototyping, laboratory qualification, and test-beam operation. The Caribou system [1, 2, 3, 4, 5, 6, 7] was developed to meet these needs within several collaborative frameworks — CERN EP R&D, DRD3, AIDAinnova, and Tangerine — and is now used by more than fourteen institutes for the characterization of a wide range of detector prototypes. It combines a custom Control and Readout (CaR) board [8] with a Xilinx Zynq System-on-Chip (SoC) platform running a common firmware and software stack shared across projects. Its architecture emphasizes modularity, reusability, and ease of integration, enabling more than fifteen different detector prototypes to be tested within a unified environment. The CaR board provides key interfaces such as programmable power supplies, voltage and current references, high-speed ADCs, injection pulsers, and configurable I/O lines for detector control and readout. It connects to the SoC through FMC and to detector-specific chip boards via a SEARAY connector, allowing adaptation to various detector designs while maintaining a common DAQ backbone. On the software and firmware side, Caribou integrates two central components via its PetaLinux image builder [9] : Peary [10], a C++ framework offering hardware abstraction, configuration management, and multi-device control via command-line and Python interfaces; and Boreal [11, 12], an FPGA firmware framework providing reusable IP cores and automated synthesis and simulation workflows. The use of Vivado Block Design facilitates the coexistence of shared control infrastructure and user-specific logic, while a strict separation between common and detector-specific components ensures maintainability and scalability. Figure 1 shows a block diagram of the Caribou system hardware architecture. It outlines the key resources provided by the Zynq SoC platform, the CaR board, and the chip board, as well as their interconnections and corresponding signal types.
Figure 1: Block diagram of the Caribou hardware architecture, illustrating all analog and digital components, as well as the system’s available connections and interfaces
@@ 2.1 Hardware Design and Validation
To validate the functionality of the CaR board v1, a dedicated validation setup was developed. Figure 2(a) shows a picture of the setup, which comprises a Zynq board, an adapter board, the CaR board under test, and a custom breakout board. The Zynq board executes the software and firmware routines used to verify the different functionalities of the CaR board. The adapter board serves as an I/O expander, providing access to all CaR board interfaces by utilizing both FMC connectors on the Zynq board. Finally, the custom breakout board enables loopback routing of LVDS signals, includes SMA connectors for high-speed transceiver lines, and provides voltage and current inputs and outputs for both the slow and fast ADCs. The primary objective of this setup is to validate all functionalities of the CaR board by testing its individual components and I/Os upon the arrival of newly produced boards and before their distribution to users. The currently available routines enable verification of the LVDS lines, high-speed transceiver links, power supplies, and ADCs, and are developed within the Boreal [11] and Peary [10] frameworks.
(a) CaR board v1 validation setup, consisting of a Zynq board, the CaR board under test, and a breakout board that provides access to all components and signals available on the CaR board
(b) Test board implementing the complete set of features selected for the v2 CaR board
Figure 2: Pictures of the CaR board v1 validation setup (a) and the v2 test board (b)
A major upgrade of the Caribou system, designated Caribou v2.0, is currently under development to transition the platform to an UltraScale+ System-on-Module architecture with a fully redesigned CaR board. While this evolution primarily targets the system-level integration and SoC platform, it also introduces substantial enhancements to the analog control and monitoring capabilities compared to the existing v1 implementation. The currently deployed CaR board (v1) will be significantly extended in the upcoming v2 revision, which is still under design. While v1 relies primarily on 12-bit DACs for power, bias, and current control, v2 will retain 12-bit resolution for most power rails but extend the available voltage ranges, in particular for higher-voltage supplies. The current source architecture will be substantially improved in v2, introducing multiple selectable bipolar ranges spanning from tens of microamperes with step sizes down to $10nA$ up to around $100mA$ with step sizes up to $50uA$ and a continuous positive/negative range, compared to the limited $1.024mA$ range of v1. Bias generation will be upgraded from a 12-bit unipolar DAC to a 16-bit multi-range DAC supporting both unipolar and bipolar outputs up to $\pm 15V$, and complemented by a dedicated bias source with programmable current and voltage limits. On the readout side, v2 will replace the low-speed monitoring ADC of v1 with a high-resolution 24-bit ADC optimized for precision measurements, while the $65MSps$ high-speed ADC present on v1 will be removed. Instead, v2 will integrate an FMC connector to support compatible expansion modules, such as high-speed ADC FMC cards. Finally, it will incorporate enhanced overcurrent protection implemented on an FPGA.
In preparation for the design of the CaR board v2, a simplified test board as shown in Figure 2(b) was developed. This $16.5\times 17.5\ \mathrm{cm^{2}}$ compact board serves as a platform to evaluate the different v2 design options and to validate the performance of previously described target components and circuits, all accessible and controlled through a single USB-to-I2C/SPI interface. The System-on-Module is not included at this stage. Two test boards have been produced and are fully operational, with extensive validation currently in progress. The high-voltage power supplies provide adjustable outputs from approximately +2.5V+2.5V to +18V+18V and 2.5V-2.5V to 18V-18V at up to $1A$, with one pair intended for the Device Under Test (DUT) and two pairs for powering the DACs, current sources, and bias circuitry. Initial DC-DC converter tests are largely complete and confirm operation up to 1718V17-18V to meet DAC headroom requirements, although further tuning and analysis of regulation, efficiency, and LDO response speed are still ongoing. The low-voltage supplies provide outputs from approximately $0.8V$ up to $5V$ at $2A$, and have been successfully tested down to $2V$. All tested current sources (two per board on two boards) are operational and show a gain error performance within $\pm 0.1-0.25\%$ of the full-scale range. The bias source provides a $\pm 15V$ output range (extendable with a larger supply range) and features a programmable current limit of $\pm 10.24mA$. The DAC provides a $\pm 15V$ output range with up to $55mA$ per channel, with the maximum current primarily limited by thermal considerations. One of the monitoring ADCs supports a specified input range of $\pm 10V$ and remains functional up to $\pm 20V$. Finally, the current protection circuit and associated FPGA logic were tested and operate as intended: the current can be read back, the current limit and response time are configurable, and channels can be grouped such that a fault in one channel disables all channels in the group.
@@ 2.3 Embedded Software Architecture Design Revision
Figure 3: Simplified UML diagram of the revised Peary architecture, illustrating the structure of the embedded application and the relationships among all system components
@@ 3 Conclusion
The developments presented in this paper mark a significant step forward in the evolution of the Caribou system. On the hardware side, the validation of the CaR board v1 and the design of the CaR board v2 test board have established a robust foundation for the upcoming CaR board v2. These efforts have enabled early testing and optimization of key analog and power-supply circuits, ensuring improved reliability and performance in the next generation of the platform. In parallel, the complete redesign of the Peary embedded software has introduced a modular and extensible architecture that can support multiple Zynq platforms and future hardware revisions. The introduction of the Hardware Abstraction Layer and the concept of Resources provides a flexible framework for integrating new boards and detector devices with minimal reconfiguration. Together, these advancements considerably strengthen the modularity, scalability, and maintainability of the Caribou system.